# Optimus: Accelerating Large-Scale Multi-Modal LLM Training by Bubble Exploitation

### v1

@@ 3.1 Design Decisions
Design decision 1: Colocate encoders and LLM with separate parallelism. To ensure that each GPU possesses both encoder and LLM model states, we propose assigning separate parallel plans to encoders and LLMs across all GPUs. This strategy is illustrated in Figure 5, where using parallel plan (DP=2, PP=2, TP=2) for encoders and (DP=1, PP=4, TP=2) for LLM. Each GPU retains both encoder and LLM model states, and then it becomes feasible for all GPUs to execute encoder computations during LLM bubbles. Note that colocating both the encoder and LLM states may require more GPU memory and we analyze the memory overhead in Section 4.5.
Figure 5: All GPUs both hold encoder and LLM model states when giving encoder and LLM separate parallel plans
@@
Figure 6: Solve complex dependencies in MLLM training through local scheduling and global ordering
@@ 3.2 Optimus  Overview
Optimus is a distributed training system designed for MLLM, enabling the scheduling of encoder computation within LLM bubbles to improve end-to-end training latency. To tackle challenges in Section 3.1, Optimus has two components, which are the model planner and bubble scheduler.
Model Planner. The model planner partitions encoders and the LLM backbone separately to all given GPUs (addressing Challenge 1 in §3.1). Initially, the planner determines the 3D parallelism plan ($\textit{DP}_{llm},\textit{PP}_{llm},\textit{TP}_{llm}$) for the LLM backbone based on insights in Megatron-LM[21]. Subsequently, the planner enumerates potential 3D parallelism plans ($\textit{DP}_{enc},\textit{PP}_{enc},\textit{TP}_{enc}$) for the encoders, considering the available GPU memory after the deployment of the LLM. With the model planner, each GPU holds both LLM and encoder model states, enabling encoder computation during LLM bubbles. The encoder and LLM model parallel plans are provided as input to the bubble scheduler, where Optimus selects parallel plans based on the output schedule with the shortest execution time.
Bubble Scheduler. Bubble scheduler is responsible for scheduling encoder computation into LLM bubbles. Given that the LLM training pipeline divides data into multiple microbatches, the scheduler schedules encoder computations on a per-microbatch basis and satisfies encoder-LLM data dependency at microbatch level (addressing Challenge 2 in §3.1). In addition, the scheduler breaks down encoder computation into kernel granularity, to enable the utilization of sub-millisecond bubbles (TP bubbles) during LLM training (addressing Challenge 3 in §3.1).
Optimus uses the model planner to devise parallel plans for both encoders and LLMs. Subsequently, for each encoder parallel plan, Optimus utilizes the bubble scheduler to generate a schedule and estimate the latency. Ultimately, Optimus selects the schedule with the shortest training time to schedule encoder computation into LLM bubbles. The workflow of Optimus is outlined in Algorithm 1.
@@ 4 Optimus Design
Section 4.1 describes how the model planer searches the parallel plans for the encoder, Section 4.2 details how the bubble scheduler exploits the coarse-grained and fined-grained bubbles through local scheduling, Section 4.3 discusses how the bubble scheduler handles encoder-LLM data dependencies through global ordering, Section 4.4 designs the bubble scheduling in multi-branch encoder models, and Section 4.5 analyzes the memory consumption of the bubble scheduling algorithm.
@@ 4.1 Model Planner
Searching separate parallel plans. Initially, the planner determines the 3D parallelism plan ($\textit{DP}_{llm},\textit{PP}_{llm},\textit{TP}_{llm}$) for the LLM backbone based on insights in Megatron-LM[21]. Subsequently, the planner enumerates potential 3D parallelism plans ($\textit{DP}_{enc},\textit{PP}_{enc},\textit{TP}_{enc}$), ensuring that $\textit{PP}_{enc}$ is a factor of $\textit{PP}_{llm}$ and $\textit{TP}_{enc}$ is a factor of $\textit{TP}_{llm}$. In practice, $\textit{PP}_{llm}$ can reach up to 64 and $\textit{TP}_{llm}$ up to 8 for training large language models (LLMs) [21]. Consequently, there are generally no more than 28 encoder parallel plans available, with up to 7 options for $PP_{enc}$ and 4 for $TP_{enc}$.
Colocating encoders and LLM. To guarantee that each GPU can perform encoder computations during LLM downtime, the model planner assigns both encoder and LLM model states to every GPU. As illustrated in Figure 5, all GPUs contain model states for both the encoder (depicted in green) and the LLM (shown in red). Without such colocation, many GPUs would lack the necessary encoder model states to execute encoder computations.
Prune parallel plans based on memory constraint. As we colocate the encoder and LLM stages on GPUs, we calculate the memory requirements for both encoder and LLM states based on the chosen parallelism plan, referencing memory analysis in [15]. Plans that violate GPU memory capacity are immediately pruned.
@@ 1 Function BubbleScheduler(encPlan, llmPlan):
Although LLM bubbles in different GPUs have different start times and duration, there is one common pattern of LLM bubbles as shown in Figure 8. There is one single big bubble (the sum of DP all-gather bubble and PP-warm bubble) before any LLM computation starts and one single big bubble (the sum of PP-cooldown bubble and reduce-scatter bubble) after all LLM computation finishes. And there are many small bubbles (PP bubbles and TP bubbles)[21, 26, 15] interleaved with LLM computation.
Figure 8: Bubble pattern of 3D parallelism
Design decision 2: The bubble scheduler, as described in Algorithm 2, initially engages in coarse-grained bubble exploitation by creating initial schedules that incorporate encoder computations within the bubbles positioned before and after LLM computations (line 2). However, it’s possible that these two bubbles may not allow sufficient time to complete all encoder computations, leading to some encoder computations being unscheduled within bubbles. To reduce the total training time, the bubble scheduler then executes fine-grained bubble exploitation. This involves refining the schedule by allocating encoder forward computations to the bubbles that alternate with LLM computations (line 7), followed by assigning encoder backward computations to these same bubbles (line 8). The final output of the bubble scheduler is the schedule that achieves the shortest possible runtime.
Coarse-grained bubble exploitation. For each potential data partitioning approach, the bubble scheduler initializes the schedule by scheduling encoder forward operations to occur before LLM computations and encoder backward operations to occur after LLM computations. Figure 9 illustrates the initialized schedule when there are $m=2$ encoder pipelines and the data partitioning approach is $[3,5]$, i.e., 3 microbatches is allocated to the first encoder pipeline and 5 for the second encoder pipeline.
@@
Fine-grained bubble exploitation. The OptimizeSchedule function (line 15 at Algorithm 2) refines the initial schedule through an iterative approach. Initially, the bubble scheduler employs findCritical to identify the encoder pipeline whose computation is on the critical path of the end-to-end MLLM training (line 17). Subsequently, it utilizes AssignKernels to allocate one microbatch of this encoder computation to bubbles interleaved with LLM computations (line 18). If there are sufficient bubbles available for scheduling encoder computation and encoder-LLM data dependencies are met, the bubble scheduler repeats this process. Otherwise, it returns the current optimized schedule.
@@
Figure 10: Find encoder pipeline that is on the critical path of end-to-end MLLM training (left: encoder forward on the critical path, right: encoder backward on the critical path)
When scheduling encoder computation to bubbles interleaved with LLM compute (AssignKernels at line 18), the bubble scheduler decomposes the encoder computation into kernel granularity and schedules these kernels based on the duration of the bubble. For each bubble, the bubble scheduler schedules multiple kernels while ensuring that the total execution time of these kernels is within the bubble duration. Additionally, the bubble scheduler must satisfy the encoder’s internal data dependencies. As illustrated in Figure 11, device 1 holds the first two layers of the encoder, while device 2 holds the next two layers. When scheduling kernels during the forward pass, device 2 can only utilize bubbles that occur after device 1 completes its forward pass to execute encoder computation. For the forward computation, the bubble scheduler schedules encoder computation from upstream encoder pipeline stages to downstream encoder pipeline stages. Conversely, for backward computation, the bubble scheduler schedules encoder computation in the reverse order. While each encoder layer also includes communication kernels, the scheduler ensures that these kernels are not assigned to TP bubbles that occur during LLM communication. Instead, the scheduler identifies long-duration computation kernels within the LLM layers and overlaps them with encoder communication kernels. As the LLM and encoder layers alternately perform computation and communication tasks, they make efficient use of GPU bandwidth and Streaming Multiprocessors (SMs). This design strategy helps to minimize resource contention and improves overall GPU utilization[16].
Figure 11: Scheduling encoder computation kernels needs to satisfy encoder internal dependencies
Complexity. Our bubble scheduling algorithm has low complexity. Given $n$ GPUs and the number of prime factors of $n$ is $n_{p}$, the search space of parallel plans is $C_{n_{p}+1}^{2}$. The number of microbatch partitioning is $O(N_{mb}^{m-1})$. Hence, the complexity for scheduling bubbles is $O(C_{n_{p}+1}^{2}*N_{mb}^{m}*(F+B))$. For our experimented settings, it usually takes around several minutes to calculate the optimal schedule (see §5.3.2), which is also a one-time cost.
@@ 4.3 Address Encoder-LLM dependency
Table 2: The list of symbols frequently used in the paper
@@
Figure 12: Interleaved 1F1B pipeline schedules before and after adjustment. The top figure shows the default interleaved 1F1B schedule in Megatron-LM paper [21]. The bottom figure shows the interleaved 1F1B schedule after adjustment. In both schedules, each device is assigned 2 chunks. Dark colors show the first chunk and light colors show the second chunk.
@@
Figure 13: Illustraction example of checkEncLLMDep
When dependencies are satisfied, the bubble scheduler integrates necessary peer-to-peer (P2P) communications into the training schedule between the last stage of the encoder pipeline and the first stage of the LLM pipeline. For instance, if encoder pipeline $j$ completes the forward pass for microbatch $i$, the scheduler will insert a P2P send (sending activations) at the last stage of encoder pipeline $j$ and a P2P receive (receiving activations) at the first stage of the LLM pipeline. Similarly, when the LLM pipeline completes the backward pass for microbatch $i$, the scheduler adds a P2P send (sending gradients) at the first stage of the LLM pipeline and a P2P receive (receiving gradients) at the last stage of encoder pipeline $j$. In the scenario depicted in Figure 13, the scheduler inserts 8 pairs of P2P send-receive at devices 1 and 2 to manage the dependencies between encoder pipeline 1 and the LLM pipeline, with 4 pairs allocated for forward dependencies and 4 pairs for backward dependencies. Likewise, an additional 8 pairs of P2P send-receive are inserted at devices 3 and 4 to address the dependencies between encoder pipeline 2 and the LLM pipeline.
@@ 4.4 Multi-Branch Encoder Scheduling
Figure 14: An example of model planner applying $\textit{PP}_{enc}=2$ to MLLM with two encoders
@@ 5 Evaluation
We have developed Optimus based on the open-source Megatron-LM framework [1] and evaluate Optimus on training large-scale mulitimodal LLMs.
@@ 5.1 Methodology
Testbed. We conduct our experiments in a production training cluster with thousands of NVIDIA Hopper GPUs. Each GPU has 80GB memory and 989TFLOPS computing performance. The intra-server connection is NVLink and the inter-server connection is a high-bandwidth RDMA network.
@@
Baselines. We use three open-sourced MLLM training systems with one strawman method as our baselines for comparison.
@@
$\bullet$ Megatron-LM [21]: Megatron-LM is a state-of-the-art LLM training framework that integrates 3D parallelism techniques. Megatron-LM is designed for symmetric transformer models, and we place multimodal encoders to the preprocess in the first pipeline stage to adapt to MLLM training.
@@ 5.2.1 Weak-Scaling Experiment
Table 3: Weak-scaling MLLM configurations
@@
Figure 15: Weak-scaling experiment results. (Alpa and FSDP are not shown in the figure because of OOM)
@@
Table 4: Training performance comparison with Alpa and FSDP
@@ 5.2.2 Strong-Scaling Experiment
Experiment setup. We assess the strong-scaling training performance of Optimus and Megatron-based baselines using the ViT-22B+GPT-175B model. Following [14], we progressively increase the number of GPUs used (1536, 2048, and 3172) while keeping the batch size constant at 1536.
Results. Table 5 compares training performance between Optimus and Megatron-LM based baselines with an increasing number of GPUs. Optimus reduces iteration time by up to 21.3% compared to Megatron-LM, and by up to 20.5% compared to the Megatron-LM balanced. With the increase in GPU count, Optimus exhibits a more pronounced speedup relative to baseline solutions. This enhanced performance is anticipated since the constant batch size coupled with an increased GPU count escalates the bubble ratio, enabling Optimus to allocate a larger proportion of encoder computations to LLM bubbles. It is also evident that Optimus maintains a stable MFU, whereas the baseline MFU declines when scaling to more GPUs.

### v2

@@ 3.1 Design Decisions
Design decision 1: Colocate encoders and LLM with separate parallelism. To ensure that each GPU possesses both encoder and LLM model states, we propose assigning separate parallel plans to encoders and LLMs across all GPUs. This strategy is illustrated in Figure 5, where using parallel plan (DP=2, PP=2, TP=2) for encoders and (DP=1, PP=4, TP=2) for LLM. Each GPU retains both encoder and LLM model states, and then it becomes feasible for all GPUs to execute encoder computations during LLM bubbles. Note that colocating both the encoder and LLM states may require more GPU memory, and we analyze the memory overhead in Section 4.5.
Figure 5: All GPUs both hold encoder and LLM model states when giving encoder and LLM separate parallel plans.
@@
Figure 6: Solve complex dependencies in MLLM training through local scheduling and global ordering.
@@ 3.2 Optimus  Overview
Optimus is a distributed training system designed for MLLM, enabling the scheduling of encoder computation within LLM bubbles to improve end-to-end training latency. To tackle the challenges outlined in Section 3.1, Optimus comprises two components: the model planner, which addresses Challenge 1 by ensuring all GPUs hold both encoder and LLM model states, and the bubble scheduler, which addresses Challenge 2 (complex dependencies in MLLM training) and Challenge 3 (sub-millisecond LLM bubbles).
Model Planner. The model planner partitions encoders and the LLM backbone separately to all given GPUs (addressing Challenge 1 in §3.1). It selects a 3D parallelism plan for the LLM backbone and explores possible 3D parallelism plans for the encoders, considering the available GPU memory after the deployment of the LLM. With the model planner, each GPU holds both LLM and encoder model states, enabling encoder computation during LLM bubbles. The encoder and LLM model parallel plans are provided as input to the bubble scheduler, where Optimus selects parallel plans based on the output schedule with the shortest execution time.
Bubble Scheduler. Bubble scheduler is responsible for scheduling encoder computation into LLM bubbles. Given that the LLM training pipeline divides data into multiple microbatches, the scheduler schedules encoder computations on a per-microbatch basis and satisfies encoder-LLM data dependency at the microbatch level (addressing Challenge 2 in §3.1). In addition, the scheduler breaks down encoder computation into kernel granularity, to enable the utilization of sub-millisecond bubbles (TP bubbles) during LLM training (addressing Challenge 3 in §3.1). The current design is based on the 1F1B-interleaved pipeline schedule [20], but it is not tightly coupled to it and can be adapted to support other pipeline schedule strategies (as discussed in Section 6).
Optimus uses the model planner to devise parallel plans for encoders and LLMs. Subsequently, for each encoder parallel plan, Optimus utilizes the bubble scheduler to generate a schedule and estimate the latency. The latency estimation is based on offline profiling conducted during the model planner step, where we use the offline profiled encoder execution time and the LLM pipeline training timeline. Ultimately, Optimus selects the schedule with the shortest training time to schedule encoder computation into LLM bubbles. The workflow of Optimus is outlined in Algorithm 1.
@@ 4 Optimus Design
Section 4.1 describes how the model planner searches the parallel plans for the encoder, Section 4.2 details how the bubble scheduler exploits the coarse-grained and fine-grained bubbles through local scheduling, Section 4.3 discusses how the bubble scheduler handles encoder-LLM data dependencies through global ordering, Section 4.4 designs the bubble scheduling in multi-branch encoder models, and Section 4.5 analyzes the memory consumption of the bubble scheduling algorithm.
@@ 4.1 Model Planner
The workflow of the model planner consists of searching encoders and LLM parallelism plans, colocating encoders and LLMs, ensuring memory constraints are met, and constructing separate microbatches for encoder and LLM pipelines.
Searching separate parallel plans. Initially, the planner determines the 3D parallelism plan ($\textit{DP}_{llm},\textit{PP}_{llm},\textit{TP}_{llm}$) for the LLM backbone based on insights in Megatron-LM[20]. Subsequently, the planner enumerates potential 3D parallelism plans ($\textit{DP}_{enc},\textit{PP}_{enc},\textit{TP}_{enc}$). To guarantee that multiple encoder models can be colocated with each LLM model, we enforce that $\textit{PP}_{enc}$ is a factor of $\textit{PP}_{llm}$ and $\textit{TP}_{enc}$ is a factor of $\textit{TP}_{llm}$.
Colocating encoders and LLM. To guarantee that each GPU can perform encoder computations during LLM downtime, the model planner assigns encoder and LLM model states to every GPU. As illustrated in Figure 5, all GPUs contain model states for both the encoder (depicted in green) and the LLM (shown in red). Without such colocation, many GPUs would lack the necessary encoder model states to execute encoder computations.
Prune parallel plans based on memory constraint. Since the encoder and LLM stages are colocated on GPUs, we estimate the memory requirements for encoder and LLM model states, as well as LLM activations, based on the selected parallelism plan—drawing on the memory analysis in [14]. We omit encoder activations from the estimation due to their negligible memory footprint. Any plans that exceed GPU memory capacity are pruned early.
@@ 1 Function BubbleScheduler(encPlan, llmPlan):
Although LLM bubbles in different GPUs have different start times and durations, there is one common pattern of LLM bubbles as shown in Figure 8. There is one single big bubble (the sum of DP all-gather bubble and PP-warm bubble) before any LLM computation starts, and one single big bubble (the sum of PP-cooldown bubble and reduce-scatter bubble) after all LLM computation finishes. And there are many small bubbles (PP bubbles and TP bubbles)[20, 25, 14] interleaved with LLM computation.
Figure 8: Bubble pattern of 3D parallelism.
The bubble scheduler, as described in Algorithm 2, initially engages in coarse-grained bubble exploitation by creating initial schedules that incorporate encoder computations within the bubbles positioned before and after LLM computations (line 2). However, it’s possible that these two bubbles may not allow sufficient time to complete all encoder computations, leading to some encoder computations being unscheduled within bubbles. To reduce the total training time, the bubble scheduler then executes fine-grained bubble exploitation. This involves refining the schedule by allocating encoder forward computations to the bubbles that alternate with LLM computations (line 7), followed by scheduling encoder backward computations to these same bubbles (line 8). The final output of the bubble scheduler is the schedule that achieves the shortest possible runtime.
Coarse-grained bubble exploitation. For each potential data partitioning approach, the bubble scheduler initializes the schedule by scheduling encoder forward operations to occur before LLM computations and encoder backward operations to occur after LLM computations. Figure 9 illustrates the initialized schedule when there are $m=2$ encoder pipelines and the data partitioning approach is $[3,5]$, i.e., 3 microbatches are allocated to the first encoder pipeline and 5 to the second encoder pipeline.
@@
Fine-grained bubble exploitation. The OptimizeSchedule function (line 15 at Algorithm 2) refines the initial schedule through an iterative approach. Initially, the bubble scheduler employs findCritical to identify the encoder pipeline whose computation is on the critical path of the end-to-end MLLM training (line 17). Subsequently, it utilizes ScheduleKernels to allocate one microbatch of this encoder computation to bubbles interleaved with LLM computations (line 18). If enough bubbles are available to accommodate encoder computation and encoder-LLM data dependencies are satisfied (as discussed in §4.3), the bubble scheduler continues scheduling. Otherwise, it terminates and returns the best schedule found so far.
@@
Figure 10: Find encoder pipeline that is on the critical path of end-to-end MLLM training (left: encoder forward on the critical path, right: encoder backward on the critical path).
When scheduling encoder computation to bubbles interleaved with LLM compute (ScheduleKernels at line 17), the bubble scheduler decomposes the encoder computation into kernel granularity and schedules these kernels based on the duration of the bubble. For each bubble, the bubble scheduler schedules multiple kernels while ensuring that the total execution time of these kernels is within the bubble duration. Additionally, the bubble scheduler must satisfy the encoder’s internal data dependencies. As illustrated in Figure 11, device 1 holds the first two layers of the encoder, while device 2 holds the next two layers. When scheduling kernels during the forward pass, device 2 can only utilize bubbles that occur after device 1 completes its forward pass to execute encoder computation. For the forward computation, the bubble scheduler schedules encoder computation from upstream encoder pipeline stages to downstream encoder pipeline stages. Conversely, for backward computation, the bubble scheduler schedules encoder computation in the reverse order. While each encoder layer also includes communication kernels, the scheduler ensures that these kernels are not assigned to TP bubbles that occur during LLM communication. Instead, the scheduler identifies long-duration computation kernels within the LLM layers and overlaps them with encoder communication kernels. As the LLM and encoder layers alternately perform computation and communication tasks, they make efficient use of GPU bandwidth and Streaming Multiprocessors (SMs). This design strategy helps to minimize resource contention and improves overall GPU utilization[15].
Figure 11: Scheduling encoder computation kernels needs to satisfy encoder internal dependencies.
Complexity. Our bubble scheduling algorithm has low complexity. Given $n$ GPUs and the number of prime factors of $n$ is $n_{p}$, the search space of parallel plans is $C_{n_{p}+1}^{2}$. The number of microbatch partitioning is $O(N_{mb}^{m-1})$. Hence, the complexity for scheduling bubbles is $O(C_{n_{p}+1}^{2}*N_{mb}^{m}*(F+B))$. In our experiments, it usually takes around several minutes to calculate the optimal schedule (see §5.3.2), which is also a one-time cost.
@@ 4.3 Address Encoder-LLM dependency
Table 2: The list of symbols frequently used in the paper.
@@
Figure 12: Interleaved 1F1B pipeline schedules before and after adjustment. The top figure shows the default interleaved 1F1B schedule in the Megatron-LM paper [20]. The bottom figure shows the interleaved 1F1B schedule after adjustment. In both schedules, each device is assigned 2 chunks. Dark colors show the first chunk, and light colors show the second chunk.
@@
Figure 13: Illustraction example of checkEncLLMDep.
When dependencies are satisfied, the bubble scheduler integrates necessary peer-to-peer (P2P) communications into the training schedule between the last stage of the encoder pipeline and the first stage of the LLM pipeline. For instance, if encoder pipeline $j$ completes the forward pass for microbatch $i$, the scheduler will insert a P2P send (sending activations) at the last stage of encoder pipeline $j$ and a P2P receive (receiving activations) at the first stage of the LLM pipeline. Similarly, when the LLM pipeline completes the backward pass for microbatch $i$, the scheduler adds a P2P send (sending gradients) at the first stage of the LLM pipeline and a P2P receive (receiving gradients) at the last stage of encoder pipeline $j$. In the scenario illustrated in Figure 13, where the training pipeline processes 8 microbatches, the scheduler inserts 8 pairs of P2P send-receive operations between devices 1 and 2 to manage the dependencies between encoder pipeline 1 and the LLM pipeline. These include 4 pairs for forward activation, send/receive, and 4 pairs for backward gradient send/receive. Similarly, another 8 pairs of P2P send-receive operations are inserted between devices 3 and 4 to handle the dependencies between encoder pipeline 2 and the LLM pipeline.
@@ 4.4 Multi-Branch Encoder Scheduling
Figure 14: An example of model planner applying $\textit{PP}_{enc}=2$ to MLLM with two encoders.
@@ 5 Evaluation
We have developed Optimus based on the open-source Megatron-LM framework [1] and evaluated Optimus on training large-scale multimodal LLMs.
@@ 5.1 Methodology
Testbed. We conduct our experiments in a production training cluster with thousands of NVIDIA Hopper GPUs. Each GPU has 80GB of memory and 989 TFLOPS of computing performance. The intra-server connection is NVLink, and the inter-server connection is a high-bandwidth RDMA network.
@@
Baselines. We compare Optimus against three open-source MLLM training systems and one strawman baseline. DiffusionPipe [30] and DistTrain [37] are excluded from baseline comparison; the rationale is detailed in Section 6.
@@
$\bullet$ Megatron-LM [20]: Megatron-LM is a state-of-the-art LLM training framework that integrates 3D parallelism techniques. Megatron-LM is designed for symmetric transformer models, and we place multimodal encoders in the pre-process in the first pipeline stage to adapt to MLLM training.
@@ 5.2.1 Weak-Scaling Experiment
Table 3: Weak-scaling MLLM configurations.
@@
Figure 15: Weak-scaling experiment results (Alpa and FSDP are not shown in the figure because of OOM).
@@
Table 4: Training performance comparison with Alpa and FSDP.
@@ 5.2.2 Strong-Scaling Experiment
Experiment setup. We assess the strong-scaling training performance of Optimus and Megatron-based baselines using the ViT-22B+GPT-175B model. Following Megascale [13], we progressively increase the number of GPUs used (1536, 2048, and 3172) while keeping the batch size constant at 1536.
Results. Table 5 compares training performance between Optimus and Megatron-LM based baselines with an increasing number of GPUs. Optimus reduces iteration time by up to 21.3% compared to Megatron-LM, and by up to 20.5% compared to the Megatron-LM balanced. As the GPU count increases, Optimus demonstrates a greater speedup over baseline solutions. While Optimus maintains a stable MFU, baseline MFU drops at larger scales. This is expected—keeping the batch size constant while scaling up the GPU count increases the bubble ratio, allowing Optimus to schedule more encoder computation into LLM bubbles.